Circular No. 6983

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1998di S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, communicate: “A spectrum of SN 1998di (cf. IAUC 6982) taken by P. Berlind on Aug. 4.4 UT at the Whipple Observatory 1.5-m telescope (+ FAST spectrograph) exhibits a blue continuum with shallow absorption troughs shortward of 500 nm. The object appears 4′′ east and 4′′ north of an anonymous galaxy with a recession velocity of 10 750 ± 30 km/s. The shallow absorption features have widths of ∼ 3000 km/s (FWHM) and correspond to He I absorption at zero radial velocity. The spectrum resembles that of CR Boo, a hydrogen-deficient dwarf nova, near maximum (Patterson et al. 1997, PASP 109, 1100). We conclude that 1998di is not a supernova, but rather is a Galactic dwarf nova discovered in outburst that is simply situated nearly in front of the distant galaxy.”

2S 1711–339 R. Remillard reports for the RXTE ASM team at Massachusetts Insti-

tute of Technology and Goddard Space Flight Center, NASA: “The RXTE All-Sky Monitor has detected x-ray emission at a position consistent with the recurrent transient 2S 1711–339 (e.g., Greenhill et al. 1979, Nature 279, 620). The ASM error circle (4′ radius) is centered 1′ north of the more accurate SAS position at α = 17^{h}14^{m}19^{s}, δ = −34^{o}02′.9 (equinox 2000.0). Retrospective analysis indicates that the x-ray flux (2–12 keV) began to increase on July 21, reaching 25 ± 4 mCrab on July 22 and 54 ± 5 mCrab on July 27. Thereafter, the ASM results are consistent with a slow decline to 33 ± 6 mCrab on Aug. 4. Radio and optical observations are needed.”

COMET C/1998 J1 (SOHO) Visual m_{1} estimates: July 3.30 UT, 8.0 (S. T. Rae, Te Awamutu, New

Zealand, 10×50 binoculars); 22.39, 9.3 (D. A. J. Seargent, Cowra, N.S.W., 15×80 binoculars); 23.42, 9.4 (Seargent).

COMET 21P/GIACOBINI-ZINNER Visual m_{1} estimates: May 31.93 UT, 14.9 (M. Lehký, Hradec Králové,

Czech Republic, 0.42-m reflector); June 25.99, 14.3 (K. Sárneczky, Ágasvár, Hungary, 0.44-m reflector); July 22.96, 13.8 (R. J. Bouma, Lattrop, The Netherlands, 0.40-m reflector); Aug. 1.97, 13.1 (M. Reszelski, Pamiatkowo, Poland, 0.25-m reflector).

1998 August 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green